You want to figure out how much ice cream can fit in the cone to the right.

To do so, you have measured that the slant height (s) is 13 cm and that the diameter (d) is 10 cm.

(Hint: you may need to calculate the height first)


1
Expert's answer
2020-11-08T18:48:01-0500

V=1/3*π*r2*h

d=2r

10=2r

r=5


132-52=122

V=1/3*π*25*12=100π ≈300